# TreeReview: A Dynamic Tree of Questions Framework for Deep and Efficient LLM-based Scientific Peer Review

### v1

@@ Abstract
Yuan Chang1,2 *thanks: * Equal contribution., Ziyue Li1,2 *, Hengyuan Zhang3 † thanks: † Corresponding author., Yuanbo Kong1,2, Yanru Wu3, Zhijiang Guo4 †, Ngai Wong5 † 1National Science Library, Chinese Academy of Sciences 2Department of Information Resources Management, School of Economics and Management, University of Chinese Academy of Sciences 3Tsinghua University 4Hong Kong University of Science and Technology (Guangzhou) 5The University of Hong Kong {changyuan, liziyue}@mail.las.ac.cn    zhang-hy22@mails.tsinghua.edu.cn
@@ 1 Introduction
The exponential growth in academic publications has placed increasing strain on the peer review system, which remains the primary quality control mechanism for scientific research Larsen and Von Ins (2010); Gropp et al. (2017). The widening gap between submission volume and reviewer availability has led to bottlenecks that potentially delay scientific progress Leopold (2015). Thus, there is an urgent need for automated methods to support the peer review process, which can provide assistance to reviewers and help authors improve their manuscripts, maintaining the quality and efficiency of scholarly evaluation Lin et al. (2023).
@@ 2 Related Work
2.1 LLMs for Review Generation
LLMs have showcased remarkable performance in several application scenarios, such as reasoning (Li et al., 2025b; Yu et al., 2025), multilingualism (Huang et al., 2023; Gurgurov et al., 2024; Zhang et al., 2024a), and text generation in specialized contexts (Yang et al., 2024a, c; Liang et al., 2024b; Tan et al., 2024; Zhang et al., 2024b; Chang et al., 2024; Li et al., 2024d). Leveraging LLMs to assist peer review has recently become an emerging research direction, exploring how LLMs can potentially augment the challenging task of scholarly evaluation Zhuang et al. (2025). Several studies have evaluated or benchmarked the performance of state-of-the-art LLMs in generating paper reviews Liang et al. (2024a); Zhou et al. (2024b, a); Du et al. (2024); Mahmoud et al. (2024). These works demonstrate that while LLMs can provide meaningful feedback, they often struggle with critical analysis and tend to produce reviews that lack the depth and specificity found in human-written reviews Liang et al. (2024a); Zhou et al. (2024b); Du et al. (2024).
@@ 3.2 Top-down Question Generation Stage
Question Generator Action Principles.$M_{q}$ is implemented as an LLM-based agent guided by a carefully crafted prompt that emphasizes the following principles:
@@ 4.1 Experimental Settings
Tasks.  We evaluate our proposed framework on two distinct review scenarios: (1) Full Review Generation: This task involves producing a comprehensive review, including summary, strengths, weaknesses, questions and ratings, mirroring the complete review process of typical academic venues; (2) Actionable Feedback Comments Generation: This task focuses on generating a list of specific, critical feedback points targeting substantive weaknesses and improvement areas in a paper. We leverage these settings to test TreeReview in handling both holistic assessments and targeted critiques.
@@ 4.4 Human Evaluation
Fig. 5 shows TreeReview consistently outperforming baselines on human evaluation across two review tasks. For full reviews, TreeReview achieves win-rates between 66.25% (against DGE) and 90.00% (against Reviewer2). For feedback comments, TreeReview surpasses the strong MARG baseline (57.50% vs. 42.50%). The high inter-evaluator agreement (overall agreement of 0.75 and Cohen $\kappa$ of 0.70) indicates reliable human judgments. These results demonstrate that our framework generates reviews and comments that better align with expert preferences compared to baseline methods.
@@ 4.5 Ablation Study
To evaluate the contributions of key components in TreeReview, we conduct ablation experiments on two variants: 1) TreeReview--dec, which removes the question tree decomposition and answer aggregation, reducing the framework to direct prompting of the LLM for the review tasks; 2) TreeReview--exp, which removes the dynamic question expansion mechanism, restricting the framework to the initial question tree.
Results in Tables 1 and 2 show that TreeReview--dec significantly degrades both quantitative accuracy and qualitative review quality, highlighting the critical role of the divide-and-conquer reasoning. For TreeReview--exp, rating prediction remains relatively robust (with minor increases in MAE/MSE), but the quality of feedback comments drops notably, with reduced ability to identify critical issues (Jaccard score decreasing from 15.33% to 12.98%) and less specific comments (ITF-IDF dropping from 4.62 to 4.01). Statistical analysis reveals that the dynamic expansion mechanism triggers expansion for 38.54% of non-leaf questions on average, generating 25.6 additional questions per review, enabling deeper probing of ambiguous or underexplored areas. Further insights from our case study (Appendix D) demonstrate that many highly aligned and specific comments stem directly from fine-grained, dynamically expanded questions.
@@
Table 3: Statistics of per-paper average token usage.
@@ Conclusion
Domain Expansion  While TreeReview has demonstrated promising effectiveness in automating paper reviews within the specific domains of deep learning and neural computation, its generalizability to other scientific disciplines remains unexplored. We believe the evidence-based answering mechanism that grounds responses in the paper’s content could potentially enable TreeReview to extend beyond its current domain. It would be exhilarating to investigate whether it can yield high-quality feedback when applied to other domains.
Multimodal Input Consideration  In this work, we do not incorporate figures as model input, as we believe that their corresponding captions and analysis within the paper already provide the essential information needed for TreeReview to generate high-quality feedback. However, given the rapid advancement of multimodal models, their potential merits deserve attention, and we plan to evaluate in future work whether incorporating such models would provide substantial benefits that outweigh their computational costs.
@@ A.2 Intermediate Question Processing
The procedure for resolving intermediate questions encompasses both answer synthesis and dynamic question expansion (§3.3). For each intermediate question $q_{i}^{\text{inter}}$ (i.e., non-leaf and non-root), the Answer Synthesizer $M_{a}$ receives the question itself and all its current sub-question-answer pairs ${(q_{i,j},a_{i,j})}_{j=1}^{\bar{n}_{i}}$. The core task is to determine if the collective evidence provided by these sub-answers is sufficient to comprehensively address $q_{i}^{\text{inter}}$. We employ Chain-of-Thought (CoT) prompting to guide $M_{a}$ through this decision process and the subsequent action. Specifically, we explicitly instruct $M_{a}$ to output its reasoning steps before generating the final output. When $M_{a}$ determines that the available information is sufficient, it synthesizes an answer by integrating and abstracting insights from the sub-answers. Otherwise, $M_{a}$ generates follow-up questions targeting the gaps. The detailed prompts are provided in the Appendix E.
@@ B.2 Baselines Implementation Details
Reviewer2Reviewer2 is a two-stage review generation framework designed to enhance the coverage and specificity of generated reviews. It consists of two fine-tuned LLMs: the first model $M_{p}$ generates aspect prompt based on the paper, and the second model $M_{r}$ produces the final review based on the paper and the aspect prompt. To facilitate training, Reviewer2 introduced a Prompt Generation with Evaluation (PGE) pipeline to annotate existing review datasets with corresponding aspect prompts.
@@ C.3 Specificity Evaluation
[…]
@@ C.5 LLM-based Alignment
where $C_{gen}$ and $C_{real}$ represent the sets of generated and reference comments, respectively. The directional intersection operators $\overrightarrow{\cap}$ and $\overleftarrow{\cap}$ represent the set of aligned elements in the right or left operand, and the intersection is defined as $\frac{|C_{gen}\overrightarrow{\cap}C_{real}|+|C_{gen}\overleftarrow{\cap}C_{real}|}{2}$.
@@ Appendix D Case Study
Appendix E Prompts used in TreeReview

### v2

@@ Abstract
Yuan Chang1,2 *thanks: * Equal contribution., Ziyue Li1,2 *, Hengyuan Zhang3 † thanks: † Corresponding author., Yuanbo Kong1,2, Yanru Wu4, Hayden Kwok-Hay So3, Zhijiang Guo5, Liya Zhu1,2 †, Ngai Wong3 1National Science Library, Chinese Academy of Sciences 2Department of Information Resources Management, School of Economics and Management, University of Chinese Academy of Sciences 3The University of Hong Kong 4Tsinghua University 5Hong Kong University of Science and Technology (Guangzhou) {changyuan, liziyue, zhuly}@mail.las.ac.cn    hengyuan.zhang88@gmail.com
@@ 1 Introduction
The exponential growth in academic publications has placed increasing strain on the peer review system, which remains the primary quality control mechanism for scientific research Larsen and Von Ins (2010); Gropp et al. (2017). The widening gap between submission volume and reviewer availability has led to bottlenecks that potentially delay scientific progress Leopold (2015). Thus, there is an urgent need for automated methods to support the peer review process, which can provide assistance to reviewers and help authors improve their manuscripts, maintaining the quality and efficiency of scholarly evaluation Lin et al. (2023a).
@@ 2 Related Work
2.1 Automated Scientific Peer Review
Automated peer review has long been explored to address the increasing burden on traditional review processes. Early efforts focused on specific aspects such as reviewer assignment Kobren et al. (2019), plagiarism detection Foltýnek et al. (2019), and paper rating recommendation Kang et al. (2018); Deng et al. (2020). More recent work has attempted to generate free-form paper reviews using small language models with supervised fine-tuning approaches Yuan et al. (2022); Lin et al. (2023b), but these models’ limited context length and comprehension capabilities make it challenging to generate nuanced and comprehensive reviews for full-length academic papers.
Recently, LLMs have showcased remarkable performance in several application scenarios, such as reasoning (Li et al., 2025b; Yu et al., 2025), multilingualism (Huang et al., 2023; Gurgurov et al., 2024; Zhang et al., 2024a), and text generation in specialized contexts (Yang et al., 2024a, c; Liang et al., 2024b; Tan et al., 2024; Zhang et al., 2024b; Chang et al., 2024; Li et al., 2024d). Leveraging LLMs to assist peer review has recently become an emerging research direction, exploring how LLMs can potentially augment the challenging task of scholarly evaluation Zhuang et al. (2025). Several studies have evaluated or benchmarked the performance of state-of-the-art LLMs in generating paper reviews Liang et al. (2024a); Zhou et al. (2024b, a); Du et al. (2024); Mahmoud et al. (2024). These works demonstrate that while LLMs can provide meaningful feedback, they often struggle with critical analysis and tend to produce reviews that lack the depth and specificity found in human-written reviews Liang et al. (2024a); Zhou et al. (2024b); Du et al. (2024).
@@ 3.2 Top-down Question Generation Stage
[…]
@@ 4.1 Experimental Settings
Tasks.  We evaluate our proposed framework on two distinct review scenarios: (1) Full Review Generation: This task involves producing a comprehensive review, including summary, strengths, weaknesses, questions and ratings, mirroring the complete review process of typical academic venues; (2) Actionable Feedback Comments Generation: This task focuses on generating a list of specific, critical feedback points targeting substantive weaknesses and improvement areas in a paper. Such actionable comments are highly valued in real-world peer review for directly helping authors identify and address major flaws, yet they pose unique challenges distinct from holistic review generation, as models must pinpoint and articulate concrete issues rather than summarizing general impressions. We leverage these settings to test TreeReview in handling both holistic assessments and targeted critiques.
@@ 4.4 Human Evaluation
Fig. 5 shows TreeReview consistently outperforming baselines on human evaluation across two review tasks. For full reviews, TreeReview achieves win-rates between 66.25% (against DGE) and 90.00% (against Reviewer2). For feedback comments, TreeReview surpasses the strong MARG baseline (57.50% vs. 42.50%). The high inter-evaluator agreement (overall agreement of 0.75 and Cohen $\kappa$ of 0.70) indicates reliable human judgments. In addition, we have calculated the Spearman’s rank correlation coefficient between the "Overall Quality" scores assigned by the LLM and the aggregated win-rate scores derived from human pairwise comparisons. The results show strong consistency ($\rho\approx 0.90$, $p<0.05$), further supporting the validity of LLM-based assessment. These results demonstrate that our framework generates reviews and comments that better align with expert preferences compared to baseline methods.
4.5 Domain Generalization Analysis
Method Comp. Tech. Clari. Const. Spec. Evi. Cons. Overall Quality
DGE 6.73 6.11 7.43 7.63 7.58 6.48 9.38 7.26
SORT 7.06 6.39 7.72 8.64 7.74 6.30 9.70 7.63
TreeReview 7.64 7.83 8.38 8.80 8.62 7.85 9.58 8.31
Table 3: LLM evaluation scores of TreeReview and representative baseline methods across quality dimensions on multi-domain dataset. Abbreviations: Comp. = Comprehensiveness, Tech. = Technical Depth, Clari. = Clarity, Const. = Constructiveness, Spec. = Specificity, Evi. = Evidence Support, Cons. = Consistency.
Method LLM-based alignment Semantic similarity
Precision Recall Jaccard High Relat. % More Spec. % SN-P SN-R SN-F1
DPW 12.85 27.73 9.68 15.87 87.07 40.83 58.37 47.77
SORT 27.92 16.71 11.71 22.22 82.10 47.02 50.35 48.34
TreeReview 34.62 29.11 17.68 25.85 93.72 52.26 56.93 54.20
Table 4: LLM-based alignment and semantic similarity evaluation against ground-truth comments for actionable feedback comments generation on multi-domain dataset. Abbreviations: High Relat.= High Relatedness, More Spec.= More Specificity.
While our benchmark primarily focuses on the AI/ML domain, we extended our evaluation to papers from other scientific fields to assess the robustness and generalization of our approach. Specifically, we curate 40 papers from Nature Communications, covering biology (13), climate science (14), and quantum science (13). We employ the same evaluation metrics and representative baselines as described in §4.2 and §4.3, thereby maintaining a consistent evaluation framework.
Tables 3 and 4 present the experimental results on this multi-domain dataset. For full review generation, TreeReview achieves the highest overall quality score of 8.31, outperforming SORT (7.63) and DGE (7.26), with notable improvements in technical depth (7.83 vs. 6.39) and evidence support (7.85 vs. 6.30). For actionable feedback generation, TreeReview maintains superior performance with the highest precision (34.62%)and Jaccard similarity (17.68%).
The performance observed on this dataset closely align with our previous findings, with TreeReview consistently outperforming baselines across almost all evaluation metrics. This consistency validates that our method’s advantages stem from fundamental improvements in review generation rather than domain-specific optimizations, supporting the broader applicability of hierarchical question decomposition for automated peer review across scientific disciplines.
@@ 4.6 Ablation Study
To evaluate the contributions of key components in TreeReview, we conduct ablation experiments on two variants: 1) TreeReview-dec, which removes the question tree decomposition and answer aggregation, reducing the framework to direct prompting of the LLM for the review tasks; 2) TreeReview-exp, which removes the dynamic question expansion mechanism, restricting the framework to the initial question tree.
Results in Tables 1 and 2 show that TreeReview-dec significantly degrades both quantitative accuracy and qualitative review quality, highlighting the critical role of the divide-and-conquer reasoning. For TreeReview-exp, rating prediction remains relatively robust (with minor increases in MAE/MSE), but the quality of feedback comments drops notably, with reduced ability to identify critical issues (Jaccard score decreasing from 15.33% to 12.98%) and less specific comments (ITF-IDF dropping from 4.62 to 4.01). Statistical analysis reveals that the dynamic expansion mechanism triggers expansion for 38.54% of non-leaf questions on average, generating 25.6 additional questions per review, enabling deeper probing of ambiguous or underexplored areas. Further insights from our case study (Appendix D) demonstrate that many highly aligned and specific comments stem directly from fine-grained, dynamically expanded questions.
@@
Table 5: Statistics of per-paper average token usage.
@@ Conclusion
LLM Limitations and Risks.TreeReview relies on LLMs for both question generation and answer synthesis. The overall review quality is thus bounded by the backend LLMs’ capabilities, such as uneven knowledge coverage of highly specialized domains. Furthermore, LLMs are prone to hallucination, which could result in propagation of factual incorrect statements in some cases. Although the evidence-based answering aggregation helps mitigate this issue by grounding answers in the paper’s content, it cannot entirely eliminate the risk of producing misleading feedback.
Multimodal Input Consideration.  In this work, we do not incorporate figures as model input, as we believe that their corresponding captions and analysis within the paper already provide the essential information needed for TreeReview to generate high-quality feedback. However, given the rapid advancement of multimodal models, their potential merits deserve attention, and we plan to evaluate in future work whether incorporating such models would provide substantial benefits that outweigh their computational costs.
@@
Acknowledgements
This work was supported by the Young Backbone Talent Program of the National Science Library, Chinese Academy of Sciences (No. E5290902) and Theme-based Research Scheme (TRS) project T45-701/22-R. We would like to thank all the anonymous reviewers for their valuable comments and constructive feedback.
@@ A.2 Intermediate Question Processing
The procedure for resolving intermediate questions encompasses both answer synthesis and dynamic question expansion (§3.3). For each intermediate question $q_{i}^{\text{inter}}$ (i.e., non-leaf and non-root), the Answer Synthesizer $M_{a}$ receives the question itself and all its current sub-question-answer pairs ${(q_{i,j},a_{i,j})}_{j=1}^{\bar{n}_{i}}$. The core task is to determine if the collective evidence provided by these sub-answers is sufficient to comprehensively address $q_{i}^{\text{inter}}$. We employ Chain-of-Thought (CoT) prompting to guide $M_{a}$ through this decision process and the subsequent action. Specifically, we explicitly instruct $M_{a}$ to output its reasoning steps before generating the final output. When $M_{a}$ determines that the available information is sufficient, it synthesizes an answer by integrating and abstracting insights from the sub-answers. Otherwise, $M_{a}$ generates follow-up questions targeting the gaps. The detailed prompts are provided in the Appendix F.
@@ B.2 Baselines Implementation Details
$Reviewer2$ is a two-stage review generation framework designed to enhance the coverage and specificity of generated reviews. It consists of two fine-tuned LLMs: the first model $M_{p}$ generates aspect prompt based on the paper, and the second model $M_{r}$ produces the final review based on the paper and the aspect prompt. To facilitate training, Reviewer2 introduced a Prompt Generation with Evaluation (PGE) pipeline to annotate existing review datasets with corresponding aspect prompts.
@@ C.3 Specificity Evaluation
$\text{ITF-IDF}=\frac{1}{W}\sum_{j=1}^{W}\left(\frac{1}{m_{j}}\sum_{i=1}^{m_{j}}\log\left(\frac{m_{j}}{O_{i}^{j}}\right)\times\log\left(\frac{W}{R_{i}^{j}}\right)\right)$ (6)
@@ C.5 LLM-based Alignment
[…]
@@ Appendix D Case Study
Appendix E Further Analysis
E.1 Tree Structure Configuration Analysis
To investigate the influence of the initial tree structure on the final review quality, we conduct an ablation study focusing on the same paper detailed in our case study (§D). This analysis specifically evaluates the robustness of our framework to variations in the initial tree structure scale.
We compare the output of our standard setting against a configuration with a more constrained initial tree.
  • Original Tree: The initial tree generation allows for a maximum of 5 sub-questions at the first level ($W^{1}_{\max}=5$), with subsequent levels having maximums of 4 and 3 sub-questions respectively.
  • Compact Tree: We reduce the maximum number of sub-questions for the first level to 4 ($W^{1}_{\max}=4$), with subsequent levels having maximums of 3 and 2.
For both configurations, the maximum number of follow-up questions generated by the dynamic expansion mechanism remains constant ($W^{\text{exp}}_{\max}=2$).
Our analysis reveals that the dynamic expansion mechanism provides significant robustness, effectively compensating for potential omissions in a less extensive initial tree. For instance, in the case study paper:
  • With the Original Tree, the initial question decomposition included an inquiry about dataset statistics: “What is the distribution of features (numerical, categorical, missing values) in the datasets used…?”
  • With the Compact Tree, this specific question was absent from the initial static structure. However, the dynamic expansion mechanism subsequently generated a semantically equivalent inquiry during the review process: “What are the key characteristics (e.g., number of columns, data types, distributions, and presence of outliers) of each dataset used…?”
This result demonstrates that TreeReview is robust to variations in the initial tree configuration. The dynamic expansion mechanism ensures that critical aspects of the paper are thoroughly examined, even if they are not captured during the initial question decomposition phase. This enhances the reliability of our framework by reducing its sensitivity to the hyperparameters governing the initial tree structure.
E.2 Error Mode Analysis
To better understand the limitations of our proposed method, we conduct a qualitative analysis of samples that scored low in the LLM-based alignment evaluation. This process reveal several recurring patterns in failure cases:
  • Multimodal Content: Our current implementation does not process multimodal inputs. Consequently, it cannot capture feedback related to visual elements. For example, human reviewer comments like “Xsync notation inconsistency, especially in Figure 1” fall outside the scope of our method’s capabilities.
  • Fine-Grained Consistency: TreeReview can be challenged by subtle inconsistencies that span different sections of a paper. For instance, in one case, the ground truth feedback pointed out confusion between “Equation (4) and the distinction between $p(Q|X^{*},Z)$ and p(Q|Z)p(Q|Z).” Correctly identifying such issues requires tracking detailed information scattered across the manuscript. Although our final aggregation step is designed to synthesize a holistic review, ensuring complete consistency at this granular level remains difficult.
E.3 Performance Analysis by Paper Type
To investigate whether the model’s performance correlates with the type of document being reviewed, we categorize the papers in our benchmark into five types: Method/Algorithm, Application, Theoretical, Analysis, and Benchmark. We then analyze the relationship between document type and performance, using the Jaccard similarity of actionable feedback comments as the primary performance indicator.
Our statistical tests across the full dataset showed no significant correlation between paper type and alignment performance. However, to probe for more subtle trends, we isolate the lowest-performing 20% of papers based on their alignment scores. In this subset, we observe that Application and Theoretical papers are slightly overrepresented compared to their overall proportions in the full dataset.
We hypothesize that this may be due to the unique demands of these paper types:
  • Application papers often focus on specific practical scenarios that require diverse, domain-specific context, which can be more difficult for an LLM to align with perfectly.
  • Theoretical papers frequently contain intricate logical reasoning and abstract concepts, challenging the LLM’s ability to generate precisely aligned technical feedback.
Appendix F Prompts used in TreeReview